# Chenjie Zhong    Zhipeng Li†,*    Shangzhi Xu Xiaohu Li

### v2

@@ 5 Conclusion
In this article, we proposed a globally convergent method that can automatically determine the number of intrinsic mode functions for Variational Mode Decomposition. For narrow-banded signals, this method does not need any prior knowledge about the scope of the IMFs number, and provide non-recursively schema to determine IMFs and corresponding central frequencies simultaneously. Experiments show that, our method can provide accurate number of IMFs and corresponding center frequencies even if the center frequencies are very close. Our implementation is performant in both accuracy and speed. Moreover, rigorous proof is given for the convergence of the method, making it theoretically reliable.
@@ 6 Acknowledgement
This study is supported by the National Natural Science Foundation of China under Grant No. 61773290 and the Fundamental Research Funds for the Central Universities (22120230311).
@@ B.3 Proof of the Zero Duality Gap
Figure S2.1: The iteration for finding the supporting baseline of FFT with respect to time series $y(t)=100\sin 20\pi t$. The 1000th(a), 2000th(b), 3000th(c), 4000th(d), 5000th(e) iteration are respectively shown.
@@ Appendix C Experiment Detail
Figure S3.2: The gap-elimination course for the found spectrum supporting line with series $y(t)=100\sin 20\pi t$. (a) is the result of source spectrum subtracting found supporting baseline, and one can see that there is some additional noise-gap in the bottom of line subtracted function. (b) is the result after our KDE process to eliminate the gap at the bottom .
@@
Figure S3.4: The iteration for finding the supporting baseline of FFT with respect to time series $y(t)=10\cos(10\pi t)+20\sin(20\pi t)$. The 1000th(a), 2000th(b), 3000th(c), 4000th(d), 5000th(e) iterations are respectively shown.
Figure S3.5: The gap-elimination course for the found spectrum supporting line with series $y(t)=10\cos 10\pi t+20\sin 20\pi t$. (a) is the result of source spectrum subtracting found supporting baseline, and one can see that there is some additional noise-gap in the bottom of line subtracted function. (b) is the result after our KDE process to eliminate the gap at the bottom .
@@
Figure S3.7: The iteration for finding the supporting baseline of FFT with respect to time series (62). The 1000th(a), 2000th(b), 3000th(c) iterations are respectively shown. Note that although the spectrum of second mode, $\cos 10\pi t$, is affected by that of the first mode, $6t^{2}$, the computed supporting baseline precisely captured the trend of the lower bound so that all the modes can be filtered out.
Figure S3.8: The gap-elimination course for the found spectrum supporting line with series (62). (a) is the result of source spectrum subtracting found supporting baseline, and one can see that there is some additional noise-gap in the bottom of line subtracted function. (b) is the result after our KDE process to eliminate the gap at the bottom.
@@
Figure S3.10: The iterations for finding the supporting baseline of FFT with respect to time series $\cfrac{1}{1.2+\cos 2\pi t}+\cfrac{\cos(32\pi t+0.2\cos 64\pi t)}{1.5+\sin(2\pi t)}$ are shown. The 1000th(a), 2000th(b), 3000th(c) iterations are respectively shown.
Figure S3.11: The noise-elimination course for the found spectrum supporting line with series $\cfrac{1}{1.2+\cos 2\pi t}+\cfrac{\cos(32\pi t+0.2\cos 64\pi t)}{1.5+\sin(2\pi t)}$. (a) is the result of source spectrum subtracting found supporting baseline, and one can see that there is some additional noise-gap in the bottom of line subtracted function. (b) is the result after our KDE process to eliminate the gap at the bottom .
@@
Figure S3.13: The iteration for finding the baseline of function $y(t)=6t+\displaystyle\sum\limits_{i=1}^{10}(13-i)\cos[(20+10i)\pi t]$, The 500th(a), 1000th(b), 1500th(c) iterations are respectively shown.
Figure S3.14: The gap-elimination course for the found spectrum supporting line with series $y(t)=6t+\displaystyle\sum\limits_{i=1}^{10}(13-i)\cos[(20+10i)\pi t]$. (a) is the result of source spectrum subtracting found supporting baseline, and one can see that there is some additional noise-gap in the bottom of line subtracted function. (b) is the result after our KDE process to eliminate the gap at the bottom .
@@
Figure S3.16: Iterations for finding the supporting baseline without extending the spectrum to MLII lead from record 100 in MIT-BIH Arrhythmia Database. The 1000th(a), 2000th(b), 3000th(c) and 4000th(d) iterations are shown respectively. The solid line are the original spectrum and the dotted line are the evaluated supporting baseline during the iterations. It takes 4140 iterations to achieve the final convergence.
Figure S3.17: The decomposition result of the MLII lead from record 100 in MIT-BIH Arrhythmia Database. The first line is the original signal and the last the residual, with the 2nd to the 16th lines are the decomposed components acquired by our workflow.
Figure S3.18: Iterations for finding the supporting baseline without extending the spectrum to V4 lead from record 124 in MIT-BIH Arrhythmia Database. The 1500th(a), 3000th(b), 4500th(c) and 6000th(d) iterations are shown respectively. The solid line are the original spectrum and the dotted line are the evaluated supporting baseline during the iterations. It takes 6899 iterations to achieve the final convergence.
Figure S3.19: The decomposition result of the V4 lead from record 124 in MIT-BIH Arrhythmia Database. The first line is the original signal and the last the residual, with the 2nd to the 15th lines are the decomposed components acquired by our workflow.

### v3

@@ 5 Conclusion
In this article, we address the long-standing open challenge of automatically determining the number of intrinsic mode functions (IMFs) and their corresponding center frequencies for Variational Mode Decomposition (VMD). We first establish a geometric insight into this modal initialization problem: the number of valid IMFs is inherently equivalent to the count of disconnected intervals of the spectrum above a critical cutting curve. Based on this insight, we formulate the problem within a discrete functional framework defined by the number of isolated intervals with respect to the cutting curve. To tackle the intractable non-continuous and non-differentiable nature of this discrete functional optimization, we introduce the cutting curve as a surrogate to transform the original problem into a continuous functional optimization problem. On this basis, we propose a globally convergent method for the automatic determination of IMF number and initial center frequencies, with rigorous mathematical proofs provided for the convexity, strong duality and global convergence of the proposed framework.
For practical engineering implementation, we further develop an efficient and high-precision numerical scheme for the proposed variational framework. Specifically, we convert the original optimal variational differential equation into a matrix form, and separate the unknown variables by introducing an extended Hadamard product with compatible broadcasting rules, which reformulates the problem into a standard system of linear equations. We also balance the magnitude order of elements in the coefficient matrix to reduce its condition number, which effectively avoids numerical ill-conditioning and enables efficient, high-precision numerical solving even for long-sequence signals.
We conduct comprehensive experiments on both synthetic narrowband signals and real-world electrocardiogram (ECG) signals. Experimental results demonstrate that our method can reliably extract physically meaningful IMFs from complex signals. It is validated that the proposed algorithm can select a reasonable number of modes to avoid over-decomposition while guaranteeing high signal reconstruction accuracy. Furthermore, the decomposed modal components derived from our method retain meaningful and discriminative features of the target waveform, which is critical for subsequent signal analysis and interpretation tasks.
This work provides a theoretically rigorous and practically robust initialization paradigm for adaptive VMD, with the potential to be extended to a wide range of signal processing scenarios.
@@ 6 Acknowledgement
This study is supported by the National Natural Science Foundation of China under Grant No. 61773290 and the Fundamental Research Funds for the Central Universities (22120230311) and Tongji University Medicine-X Interdisciplinary Research Initiative (Grant No. 2025-0554-YB-11).
@@ B.3 Proof of the Zero Duality Gap
Figure S2.1: The iteration for finding the cutting curve of FFT with respect to time series $y(t)=100\sin 20\pi t$. The 1000th(a), 2000th(b), 3000th(c), 4000th(d), 5000th(e) iteration are respectively shown.
@@ Appendix C Experiment Detail
Figure S3.2: The gap-elimination course for the found spectrum supporting line with series $y(t)=100\sin 20\pi t$. (a) is the result of source spectrum subtracting found cutting curve, and one can see that there is some additional noise-gap in the bottom of line subtracted function. (b) is the result after our KDE process to eliminate the gap at the bottom .
@@
Figure S3.4: The iteration for finding the cutting curve of FFT with respect to time series $y(t)=10\cos(10\pi t)+20\sin(20\pi t)$. The 1000th(a), 2000th(b), 3000th(c), 4000th(d), 5000th(e) iterations are respectively shown.
Figure S3.5: The gap-elimination course for the found spectrum supporting line with series $y(t)=10\cos 10\pi t+20\sin 20\pi t$. (a) is the result of source spectrum subtracting found cutting curve, and one can see that there is some additional noise-gap in the bottom of line subtracted function. (b) is the result after our KDE process to eliminate the gap at the bottom .
@@
Figure S3.7: The iteration for finding the cutting curve of FFT with respect to time series (65). The 1000th(a), 2000th(b), 3000th(c) iterations are respectively shown. Note that although the spectrum of second mode, $\cos 10\pi t$, is affected by that of the first mode, $6t^{2}$, the computed cutting curve precisely captured the trend of the lower bound so that all the modes can be filtered out.
Figure S3.8: The gap-elimination course for the found spectrum supporting line with series (65). (a) is the result of source spectrum subtracting found cutting curve, and one can see that there is some additional noise-gap in the bottom of line subtracted function. (b) is the result after our KDE process to eliminate the gap at the bottom.
@@
Figure S3.10: The iterations for finding the cutting curve of FFT with respect to time series $\cfrac{1}{1.2+\cos 2\pi t}+\cfrac{\cos(32\pi t+0.2\cos 64\pi t)}{1.5+\sin(2\pi t)}$ are shown. The 1000th(a), 2000th(b), 3000th(c) iterations are respectively shown.
Figure S3.11: The noise-elimination course for the found spectrum supporting line with series $\cfrac{1}{1.2+\cos 2\pi t}+\cfrac{\cos(32\pi t+0.2\cos 64\pi t)}{1.5+\sin(2\pi t)}$. (a) is the result of source spectrum subtracting found cutting curve, and one can see that there is some additional noise-gap in the bottom of line subtracted function. (b) is the result after our KDE process to eliminate the gap at the bottom .
@@
Figure S3.13: The iteration for finding the cutting curve of function $y(t)=6t+\displaystyle\sum\limits_{i=1}^{10}(13-i)\cos[(20+10i)\pi t]$, The 500th(a), 1000th(b), 1500th(c) iterations are respectively shown.
Figure S3.14: The gap-elimination course for the found spectrum supporting line with series $y(t)=6t+\displaystyle\sum\limits_{i=1}^{10}(13-i)\cos[(20+10i)\pi t]$. (a) is the result of source spectrum subtracting found cutting curve, and one can see that there is some additional noise-gap in the bottom of line subtracted function. (b) is the result after our KDE process to eliminate the gap at the bottom .